OpenStudy (anonymous):

how would i simplify 4x-16 using -1 as a ------ factor 4-x

OpenStudy (anonymous):

did you mean (4x-16)/(4-x) ?

OpenStudy (anonymous):

-(16-4x) / (4-x) = -4(4-x)/(4-x) = -4